# Grounded GUI Understanding for Vision Based Spatial Intelligent Agent: Exemplified by Virtual Reality Apps

### v3

@@ 8. Related Work
8.1. Studies on VR/AR Apps
@@
8.2. VR/AR Testing
VR/AR (XR) apps provide an immersive experience to users, involving various procedures like device tracking and rendering. To ensure users’ experience, there are several works conducting empirical studies on XR software testing (Andrade et al., 2019). Andrade et al. (Andrade et al., 2019) compared open-source VR projects with non-VR projects to point out the necessity of performing testing on VR apps. Rzig et al. (Rzig et al., 2023) performed a large-scale empirical study on software testing practices of open-source VR projects to identify the current state of testing in VR apps. Several works have proposed approaches to facilitate XR testing. Qin and Hassan (Qin and Hassan, 2022) proposed DyTRec, which provides developers with recommendations on which codes should be tested by actually running the app. Souza et al. (Souza et al., 2018) proposed VR-ReST, a tool designed to assist with requirements specification and test data generation for VR apps. Rafi et al. (Rafi et al., 2022) proposed PredART that can be used as the test oracle when checking the placement of virtual objects in AR. Wang (Wang, 2022) proposed VRTest that extracts information from the VR scene and automatically explores the scene and interacts with the objects by controlling the camera. Andrade et al. (de Andrade et al., 2023) proposed a novel approach that combines metamorphic testing, agent-based testing, and machine learning to test VR apps. Recently, Li et al. (Li et al., 2024) proposed StereoID to automatically detect stereoscopic visual inconsistencies in VR apps.
8.3. User Interface Analysis
Empirical studies have been conducted to give insights into UI analysis (Ye et al., 2021), such as Chen et al. (Chen et al., 2020) conducted a systematic large-scale empirical study on GUI element detection methods. Previous works have adapted both old-fashioned methods  (Yeh et al., 2009; Bao et al., 2015; Dixon and Fogarty, 2010) and deep learning models (Wang et al., 2023a; Qian et al., 2022; Wu et al., 2023) to UI analysis. White et al. (White et al., 2019) replaced methods that involve GUI APIs with machine learning techniques to find interactable elements in GUI images. Xie et al. (Xie et al., 2020) proposed UIED, which combines old-fashioned computer vision approaches and deep learning models to detect components on complex GUI images. For game apps, Ye et al.  (Ye et al., 2021) conducted an empirical study of clickable GUI element detection on mobile games and constructed the first clickable GUI element detection benchmark. Wu et al. (Wu et al., 2023) constructed a larger and more precise GUI dataset for mobile games and evaluated seven GUI detection techniques on it. The industry also applies UI analysis techniques to solve business problems like GUI testing. Ran et al.  (Ran et al., 2022) reported their experiences and lessons learned developing and deploying VTest, an automated visual testing framework for smartphones that leverages only GUI images. Wang et al.  (Wang et al., 2023a) proposed iExplorerGame, a unified framework for black-box mini-game testing that combines deep learning object detection and edge aggregation-based segmentation to guide the testing. Qian et al.  (Qian et al., 2022) proposed a fast OCR-based widget localization technique Label Text Screening, which accelerates the OCR widget detection by analyzing and leveraging the feature of texts in widgets. Research on UI analysis is going in-depth and reaching fields with complex GUI-like game apps. However, existing methods are limited to common IGE with finite predefined categories, which fail to tackle the open vocabulary challenge, as mentioned in  section 1.
@@ 9. Conclusion
In this paper, we analyze the key challenges in interactable GUI element (IGE) detection for Virtual Reality (VR) apps. We propose Orienter, the first zero-shot, context-sensitive IGE detection framework for VR apps. Extensive experiments have been conducted to verify the effectiveness of the proposed IGE detection framework. The results demonstrate that Orienter is more effective than the state-of-the-art GUI element detection approaches, Experiments also illustrate that Orienter is beneficial for boosting automated GUI testing. Alongside developing Orienter, we have also created the first dedicated dataset for IGE detection in VR environments, both of which are publicly available to spur further research in this field.

### v4

@@ 8. Related Work
8.1. IGE Detection
Empirical studies have been conducted to give insights into IGE detection (Ye et al., 2021), e.g., Chen et al. (Chen et al., 2020) conducted a systematic large-scale empirical study on GUI element detection methods. Previous works have adapted both old-fashioned methods  (Yeh et al., 2009; Bao et al., 2015; Dixon and Fogarty, 2010) and deep learning models (Wang et al., 2023a; Qian et al., 2022; Wu et al., 2023) to IGE detection. White et al. (White et al., 2019) replaced methods that involve GUI APIs with machine learning techniques to find interactable elements in GUI images. Xie et al. (Xie et al., 2020) proposed UIED, which combines old-fashioned computer vision approaches and deep learning models to detect components on complex GUI images. For game apps, Ye et al.  (Ye et al., 2021) conducted an empirical study of clickable GUI element detection on mobile games and constructed the first clickable GUI element detection benchmark. Wu et al. (Wu et al., 2023) constructed a larger and more precise GUI dataset for mobile games and evaluated seven IGE detection techniques on it. The industry also applies IGE detection techniques to solve business problems like GUI testing (Ran et al., 2022; Wang et al., 2023a; Qian et al., 2022). Ran et al. (Ran et al., 2022) reported their experiences and lessons learned developing and deploying VTest, an automated visual testing framework for smartphones that leverages only GUI images. Wang et al.  (Wang et al., 2023a) proposed iExplorerGame, a unified framework for black-box mini-game testing that combines deep learning object detection and edge aggregation-based segmentation to guide the testing. Qian et al.  (Qian et al., 2022) proposed a fast OCR-based widget localization technique Label Text Screening, which accelerates the OCR widget detection by analyzing and leveraging the feature of texts in widgets. Research on UI analysis is going in-depth and reaching fields with complex GUI-like game apps. However, existing methods are limited to common IGE with finite predefined categories, which fail to tackle the open vocabulary challenge.
8.2. VR/AR Testing
VR/AR (XR) apps provide an immersive experience to users, involving various procedures like device tracking and rendering. To ensure users’ experience, there are several works conducting empirical studies on XR software testing (Andrade et al., 2019). Andrade et al. (Andrade et al., 2019) compared open-source VR projects with non-VR projects to point out the necessity of performing testing on VR apps. Rzig et al. (Rzig et al., 2023) performed a large-scale empirical study on software testing practices of open-source VR projects. Several works have proposed approaches to facilitate XR testing. Qin and Hassan (Qin and Hassan, 2022) proposed DyTRec, which provides developers with recommendations on which codes should be tested by actually running the app. Souza et al. (Souza et al., 2018) proposed VR-ReST, a tool designed to assist with requirements specification and test data generation for VR apps. Rafi et al. (Rafi et al., 2022) proposed PredART which can be used as the test oracle when checking the placement of virtual objects in AR. Wang (Wang, 2022) proposed VRTest that extracts information from the VR scene and automatically explores the scene and interacts with the objects by controlling the camera. Andrade et al. (de Andrade et al., 2023) proposed a novel approach that combines metamorphic testing, agent-based testing, and machine learning to test VR apps. Recently, Li et al. (Li et al., 2024) proposed StereoID to automatically detect stereoscopic visual inconsistencies in VR apps.
8.3. Exploratory Studies on VR/AR Apps
@@ 9. Conclusion
In this paper, we analyze the key challenges in interactable GUI element (IGE) detection for Extended Reality (XR) apps. We propose Orienter, the first zero-shot, context-sensitive IGE detection framework for XR apps. Extensive experiments have been conducted to verify the effectiveness of the proposed IGE detection framework. The results demonstrate that Orienter is more effective than the state-of-the-art GUI element detection approaches, Experiments also illustrate that Orienter is beneficial for boosting automated GUI testing. Alongside developing Orienter, we have also created the first dedicated dataset for IGE detection in XR environments, both of which are publicly available to spur further research in this field.